Big Bang Inc. produces custom model kits and uses normal costing in its job-costing system.  Selected financial information for the year just ended is as follows:

Raw Materials Control

Finished Goods Control

Beg.Bal.       $ 4,000

14,000

Beg.Bal.     $ 10,000

60,000

13,000

70,000

End.Bal.    $3,000

End.Bal.     $ 20,000

Work-in-Process Control

Beg.Bal.          $ 8,000

Direct labour   28,000

Other Information:

  • Actual manufacturing overhead was $42,000 for the year.
  • The direct labour hourly wage was $16 per hour.
  • Manufacturing overhead was allocated using an allocation rate of $20 per Direct Labour Hour.
  • Indirect materials used during the year amounted to $2,000.

Required:

  1. What was the amount of raw material purchased during the year?
  2. What was the amount of direct material issued to production (used in manufacturing) during the year?
  3. What was the amount of manufacturing overhead allocated to jobs during the year?
  4. What was the total cost of jobs completed (Cost of Goods Manufactured) during the year?
  5. What was the balance of Work in Process control at the end of the year (before disposition of over- or under allocated overhead)?
  6. What was the amount of over- or under-allocated overhead for the year?
  7. What was the cost of goods sold for the year (before disposition of over- or under allocated overhead)?
  8. Dispose of the over- or under-allocated overhead using the following:
    1. Write off to Cost of Goods Sold
    2. Proration based on ending balances (before proration) in Work-In-Process, Finished Goods Control and Cost of Goods Sold.
  9. Which method in requirement h) do you recommend? Why?
Answer #1

Answer (i):

From the T account given for Raw Material control we observe that:

Raw material purchased during the year = $13,000

Answer (ii):

From the T account given for Raw Material control we observe that:

Raw material issued = $14,000

Given:

Indirect materials used during the year amounted to $2,000.

Hence:

Direct material issued to production (used in manufacturing) during the year = 14000 - 2000 = $12,000

Direct material issued to production (used in manufacturing) during the year = $12,000

Answer (iii):

Actual Direct hours = 28000 / 16 = 1,750 DLH

Manufacturing overhead was allocated using an allocation rate of $20 per Direct Labor Hour.

Amount of manufacturing overhead allocated to jobs during the year = 1750 * $20 = $35,000

Amount of manufacturing overhead allocated to jobs during the year = $35,000

Answer (iv):

From the T account given for Finished Goods control we observe that:

Total cost of jobs completed (Cost of Goods Manufactured) during the year = $70,000

Answer (v):

Ending balance in work in process = Beginning balance in work in process + Direct material + Direct labor + Allocated overhead - Transferred cost of completed jobs

= 8000 + 12000 + 28000 + 35000 - 70000

= $13,000

Balance of Work in Process control at the end of the year = $13,000

Answer (vi):

Amount of under-allocated overhead for the year = 42000 - 35000 = $7,000

Amount of under-allocated overhead for the year = $7,000

Answer (vii):

From the T account given for Finished Goods control we observe that:

Cost of goods sold for the year (before disposition of over- or under allocated overhead) = $60,000

As HOMEWORKLIB's policy 4 parts need to be answered; I have answered 7 parts